OpenStudy (anonymous):

Which equation is correct for the perpendicular bisector of the line segment whose endpoints are (4,-3) and (8,3)? 1.) y-3= -2/3(x-6) 2.) y+3= 2/3(x-6) 3.) y= 2/3(x+6) 4.) y= -2/3(x-6)

OpenStudy (shamil98):

take the slope of the two points. and input into the formula: y - y1 = m(x-x1) where (4,-3) -->(x1,y1) and m = slope

OpenStudy (anonymous):

I got to the ACTUAL equation which is y=-2/3x+4. I just have no idea about these choices.

OpenStudy (anonymous):

Uhhh.. hold up.

OpenStudy (anonymous):

The actual slope or the reciprocal slope?

OpenStudy (shamil98):

actual slope

OpenStudy (anonymous):

The actual slope is 3/2

OpenStudy (anonymous):

The reciprocal is -2/3

OpenStudy (anonymous):

Yeah that doesn't make much sense. When I did what you said I got: y--3= -2/3 (x-4)

OpenStudy (nikato):

then u want to find the midpoint of those 2 points

OpenStudy (anonymous):

The uhh midpoint is (6,0)

OpenStudy (nikato):

so now find the equation with the new slope and that point^

OpenStudy (anonymous):

I did the SRME method. S(slope)= 3/2 R(reciprocal)= -2/3 M(midpoint)= (6,0) E(equation)= y=-2/3x+4

OpenStudy (nikato):

how did u get that equation? i think u did something wrong. i actually got one of ur choices

OpenStudy (anonymous):

Uhmm because I did: y=-2/3x+b 0=-2/3(6)+b 0=-4+b b=4 So then plug it all in together: y=-2/3x+4

OpenStudy (anonymous):

But its not one of the choices

OpenStudy (nikato):

try plugging it in in the point slope equation

OpenStudy (anonymous):

The what?

OpenStudy (nikato):

point slope equation

OpenStudy (nikato):